# Can Large Language Models Generate High-quality Patent Claims?

### v1

@@ 5 Conclusion
We construct a dataset for description-based patent claim generation and evaluate five LLMs. Our results indicate that description-based claim generation surpasses previous abstract-based methods. We also observe that current patent-specific LLMs exhibit inadequate performance, underscoring the need for further research. Moreover, while LLMs can output high-quality first independent claims, their effectiveness significantly diminishes with subsequent dependent claims. Additionally, fine-tuning can enhance feature completeness, conceptual clarity, and feature linkage, whereas multi-task fine-tuning reduces conceptual clarity. GPT-4 outperforms other models across all human evaluation metrics, demonstrating better feature coverage, conceptual clarity, and technical coherence. Despite promising, the models are not yet satisfactory for practical applications.
@@
We acknowledge several limitations of our research. Firstly, we contribute pioneering work on patent claim generation, focusing on patent descriptions limited to fewer than 8,000 tokens. Our results show that LLMs generally underperform with documents of such length. Since longer documents may pose more challenges for LLMs, our conclusions are not compromised. Secondly, our investigation is restricted to English-language patents. Future work may focus on patents in other languages. Thirdly, the effects of hyperparameter tuning are not explored in this research. More appropriate hyperparameters for training and inference may improve the results. Fourthly, the scope of human evaluation for generated patent claims is limited because of the prohibitively high costs. A promising research direction is to develop automated evaluation methods with better alignment with human evaluations in the patent domain.
@@
Appendix A Examples of Patent Description, Abstract, and Claims
Table 9 shows an example of patent description, abstract, and claims.
Appendix B Experimental Details
All fine-tuning and inference processes are conducted on NVIDIA A100 GPUs. The total running time is about 700 hours. We use the first 90% documents of the entire train set as the training dataset and the remaining 10% as the validation set. During fine-tuning, we use a LoRA rank of 96, LoRA alpha of 32, dropout of 0.05, batch size of 1, learning rate of 5e-5, weight decay of 0.1, and training epochs of 5. The training will stop early if the loss does not decrease after 500 steps. For inference, we set the temperature to 0.1 and the maximum generation tokens to 1,024. For automated evaluation metrics, we use the package from the HuggingFace Evaluate library. [^7] https://github.com/huggingface/evaluate
Appendix C Human Rating Strategies
Appendix D More Results
D.1 Statistics of Generated Claims
Gold standard 962 11.36 84.69 1.01 31.96
GPT-4 482 6.59 73.05 1.05 16.52
Llama-3 990 19.20 51.55 0.62 21.57
Llama-3-FT 836 14.49 57.70 1.09 30.40
Llama-3-MFT 857 15.36 55.81 1.11 29.66
D.2 Evaluation on Abstracts
Model BLEU R-1 R-L BERTScore
Patent-GPT-J 7.71 22.46 18.07 79.23
GPT-4 15.11 46.78 30.95 87.61
Llama-3 24.09 50.93 40.08 88.65
Llama-3-FT 48.48 67.90 59.66 92.13
Llama-3-MFT 48.49 67.69 59.65 92.13
Table 8: Evaluation results of abstract
@@ D.2 Evaluation on Abstracts
Embodiments of the present invention will be described in detail below. In the present specification, the same or equivalent components are designated by the same reference numerals. First Embodiment FIG. 1A is a cross-sectional view illustrating a structure of a semiconductor light-emitting element (hereinafter, may be referred to simply as a light-emitting element or an element) 10 according to a first embodiment. The semiconductor light-emitting element 10 has a structure in which a semiconductor structure layer SS is formed on a mounting substrate (hereinafter, may be referred to simply as a substrate) 11 formed from sapphire. The semiconductor structure layer SS will be concretely described below.

### v2

@@ 5 Conclusion
We construct a dataset for description-based patent claim generation and evaluate various LLMs. Our results indicate that description-based claim generation surpasses previous abstract-based methods. We also observe that current patent-specific or legal-specific LLMs exhibit inadequate performance, underscoring the need for further research. Moreover, while LLMs can output high-quality first independent claims, their effectiveness significantly diminishes with subsequent dependent claims. Additionally, fine-tuning can enhance feature completeness, conceptual clarity, and feature linkage, whereas multi-task fine-tuning reduces conceptual clarity. GPT-4 outperforms other models across all human evaluation metrics, demonstrating better feature coverage, conceptual clarity, and technical coherence. Despite promising, the models are not yet satisfactory for practical applications.
@@
We acknowledge several limitations of our research. Firstly, we contribute pioneering work on patent claim generation, focusing on patent descriptions limited to fewer than 8,000 tokens. Our results show that LLMs generally underperform with documents of such length. Since longer documents may pose more challenges for LLMs, our conclusions are not compromised. Secondly, our investigation is restricted to English-language patents. Future work may focus on patents in other languages. Thirdly, the effects of hyperparameter tuning are not explored in this research. More appropriate hyperparameters for training and inference may improve the results.
@@
Appendix A Patent Description, Abstract, and Claims
@@
Embodiments of the present invention will be described in detail below. In the present specification, the same or equivalent components are designated by the same reference numerals. First Embodiment FIG. 1A is a cross-sectional view illustrating a structure of a semiconductor light-emitting element (hereinafter, may be referred to simply as a light-emitting element or an element) 10 according to a first embodiment. The semiconductor light-emitting element 10 has a structure in which a semiconductor structure layer SS is formed on a mounting substrate (hereinafter, may be referred to simply as a substrate) 11 formed from sapphire. The semiconductor structure layer SS will be concretely described below. An n-type semiconductor layer 12 serving as a first semiconductor layer is formed from, for example, a GaN layer containing an n-type dopant (for example, Si).
@@
Table 7 shows an example of patent description, abstract, and claims. The patent description provides a comprehensive illustration of the invention, including structural and functional properties of the invention based on detailed embodiments and examples as well as the context in which it is used. The abstract is a short summary that provides a quick overview of the technical invention in the patent document. It is usually limited to a few sentences and must be concise, covering the main technical aspects without going into details. Patent claims are the legally significant part of a patent application, defining the technical scope of the invention to withstand legal scrutiny and secure legal protection. Claims must be crafted with high-level precision and clarity, comprising technological aspects that warrant protection and include the features that constitute the invention.
Gold standard 962 11.36 84.69 1.01 31.96
GPT-4 482 6.59 73.05 1.05 16.52
Llama-3-8B 990 19.20 51.55 0.62 21.57
Llama-3-FT 836 14.49 57.70 1.09 30.40
Llama-3-MFT 857 15.36 55.81 1.11 29.66
Llama-3-70B 760 11.15 68.15 1.62 15.29
Mistral-7B 868 17.06 50.89 0.90 26.84
Mixtral-8$\times$7B 969 13.00 74.51 0.74 25.90
Mixtral-8$\times$22B 732 11.68 62.73 0.84 28.19
Appendix B Model Details
GPT-4 We use the latest GPT-4 model (OpenAI, 2023), representing the state-of-the-art LLM with substantial general knowledge and outstanding reasoning capabilities.888gpt-4-turbo-2024-04-09: https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo Notably, GPT-4 shows remarkable performance on summarization tasks (Pu et al., 2023). Hence, we do not evaluate other LLMs specifically trained for summarization.
Llama-3 Among open-source LLMs, we choose the recent Llama-3, pre-trained on over 15 trillion tokens of data from publicly available sources. [^9] https://llama.meta.com/llama3/ Llama-3 outperforms most open-source models on common industry benchmarks. We opt for both the Llama-3-8B-Instruct and Llama-3-70B-Instruct versions because they have been fine-tuned to execute user instructions more accurately and effectively.
Llama-3-FT Llama-3-FT is a fine-tuned version of the original Llama-3-8B-Instruct model based on our train set. We use LoRA (Hu et al., 2021) for training, a parameter-efficient approach to reduce computational needs while maintaining comparable performance. The inputs are instruction prompts and patent descriptions, while the output is corresponding patent claims. Appendix C lists experimental details.
Llama-3-MFT Multi-task fine-tuning (MFT) has shown effectiveness in many applications, such as code generation (Liu et al., 2023a). MFT trains the model on multiple tasks simultaneously to enhance its ability to better generalize across a broader range of tasks by learning shared and task-specific features. Our Llama-3-MFT model is trained on two tasks, description-based claims generation and summarization (abstract generation). Both tasks require processing complex documents and generating shorter texts. Thus, we aim to investigate whether MFT could bring extra benefits to claim generation compared to single fine-tuning.
Mistral Mistral is an open-sourced LLM designed for superior performance and efficiency, which outperforms Llama-2-13B on some tasks (Jiang et al., 2023). We use the version of Mistral-7B-Instruct-v0.3 that has been instruction-tuned to follow users’ prompts. [^10] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3
SaulLM SaulLM-7B is designed for the legal domain, which adopts the Mistral-7B architecture and is trained on an English legal corpus of over 30 billion tokens (Colombo et al., 2024). The dataset includes 4.7B tokens of patent documents from the United States Patent and Trademark Office (USPTO). We use the SaulLM-7B-Instruct version for experiments.
Mixtral Mixtral-8$\times$7B is based on Sparse Mixture of Experts (SMoE) (Jiang et al., 2024). It has the same architecture as Mistral-7B, but each layer comprises 8 feedforward blocks (i.e. experts). A router network in each layer selects two experts to process the current state and combine the outputs. This model only uses 13B active parameters during inference, but outperforms the 70B version of Llama-2. We evaluate both Mixtral-8$\times$7B-Instruct and Mixtral-8$\times$22B-Instruct.
Appendix C Experimental Details
All fine-tuning and inference processes are conducted on NVIDIA A100 GPUs. The total running time is about 700 hours. We use the first 90% documents of the entire train set as the training dataset and the remaining 10% as the validation set. During fine-tuning, we use a LoRA rank of 96, LoRA alpha of 32, dropout of 0.05, batch size of 1, learning rate of 5e-5, weight decay of 0.1, and training epochs of 5. The training will stop early if the loss does not decrease after 500 steps. For inference, we set the temperature to 0.1 and the maximum generation tokens to 1,024. We have employed a standard prompt format to maintain consistency and focus on assessing the capabilities and limitations of different LLMs in this domain. The prompt is: You are a patent expert. Given the following patent description, write patent claims. For automated evaluation metrics, we use the package from the HuggingFace Evaluate library. [^11] https://github.com/huggingface/evaluate
Appendix D More Results
Metric Mistral-7B Llama-3-8B Llama-3-FT Llama-3-MFT Llama-3-70B Mixtral-8$\times$7B Mixtral-8$\times$22B GPT-4
BLEU
H (35%) 31.83 36.18 38.50 37.13 39.46 43.19 40.74 17.56
G (25%) 27.62 34.60 35.41 36.46 30.02 27.37 29.85 13.64
B (14%) 31.16 33.82 43.03 41.71 41.21 43.76 38.49 19.73
A (12%) 24.74 31.76 31.39 31.96 25.87 21.04 18.08 10.91
Others (14%) 30.77 32.88 39.07 39.73 38.59 36.86 33.73 15.85
R-1
H (35%) 50.69 58.20 62.09 59.86 60.31 64.10 62.79 54.42
G (25%) 47.13 56.88 58.33 57.79 59.03 57.09 58.77 51.03
B (14%) 49.66 57.09 63.80 62.11 62.48 64.90 62.50 56.52
A (12%) 46.44 54.47 53.88 53.91 53.26 48.80 51.27 46.34
Others (14%) 50.83 56.36 59.19 59.07 61.75 61.05 58.97 52.07
R-L
H (35%) 37.54 39.37 49.44 47.74 42.84 46.91 48.89 40.23
G (25%) 35.16 38.73 46.46 45.65 40.17 41.01 45.51 36.40
B (14%) 37.45 38.72 50.65 49.29 43.88 48.16 48.29 40.28
A (12%) 32.08 35.65 39.90 39.00 33.75 32.39 33.95 31.23
Others (14%) 36.81 37.97 46.87 46.52 44.16 42.75 45.98 36.71
BERTScore
H (35%) 85.73 88.74 90.00 89.83 87.83 89.16 89.59 87.78
G (25%) 84.79 88.27 89.16 89.17 86.87 88.17 88.85 86.88
B (14%) 85.81 88.97 90.43 90.18 88.09 89.42 89.19 87.94
A (12%) 84.35 87.50 87.42 87.85 85.59 85.54 86.49 86.29
Others (14%) 85.59 88.56 89.34 89.32 87.95 88.49 88.97 87.22
Table 9: Automated evaluation results of each patent category at the International Patent Classification (IPC) section level. Four of the most frequent sections are selected (H: Electricity. G: Physics. B: Performing operations; Transporting. A: Human necessities). The scores of the same model vary significantly among sections.
D.1 Statistics of Generated Claims
Patent texts have unique requirements to pass rigorous patent scrutiny. For example, the precision requirement and information density of patent texts are higher than in everyday language. The patent language focuses more on precision and accuracy instead of readability. Such precision requirement typically leads to high repetitiveness in both terminology and structure of sentences, paragraphs, and sections. Furthermore, sentences are often overburdened because they use relative or adverbial clauses to include specifications for precision or add examples for a wider scope. LLMs, such as GPT-4, which is trained predominantly on more accessible texts, naturally tend towards generating more readable content. Thus, this difference lets LLMs output more readable text than the gold claims, but it is irrelevant to the human expert assessment, which aims at formal and legal aspects.
D.2 Evaluation on Abstracts
Model BLEU R-1 R-L BERTScore
Patent-GPT-J 7.71 22.46 18.07 79.23
GPT-4 15.11 46.78 30.95 87.61
Llama-3 24.09 50.93 40.08 88.65
Llama-3-FT 48.48 67.90 59.66 92.13
Llama-3-MFT 48.49 67.69 59.65 92.13
Table 10: Evaluation results of abstract
D.3 Evaluation on IPC Labels
Normal Prompt Alternative Prompt
Model BLEU R-1 R-L BERTScore BLEU R-1 R-L BERTScore
SaulLM-7B 12.86 30.68 23.07 80.24 12.99 31.45 23.71 80.41
Mistral-7B 29.70 49.17 36.20 85.33 37.87 59.72 43.49 87.81
Llama-3-8B 35.42 57.17 38.88 88.49 35.61 57.50 39.17 88.51
Llama-3-70B 36.40 59.89 41.79 87.44 37.71 60.43 41.89 87.86
Mixtral-8$\times$7B 37.03 60.18 43.67 88.51 37.86 60.04 43.25 88.53
GPT-4 15.73 52.59 37.73 87.34 28.39 59.53 40.61 87.72
Table 11: Automated evaluation results on different prompts.
We report the automated evaluation results of each patent category at the International Patent Classification (IPC) section level in Table 9. We select four of the most frequent sections, including H (35%): Electricity; G (25%): Physics; B (14%): Performing operations, Transporting; A (12%): Human necessities; and others (14%). Table 9 reveals that all models exhibit fluctuations in performance across different IPC sections. For example, the performance of each metric of the same model in section B is consistently higher than others. In contrast, models usually perform worse on patents of section A. The fluctuations can be more than 15%, for instance, Mistral-8$\times$7B achieves 48.16% R-L in section B, but only 32.39% in section A. The section-wise variance suggests that certain sections may be inherently more challenging for models to handle, possibly due to the technical complexity or uniqueness of the language used in those sections. Overall, the results highlight the challenges in achieving uniformly high performance across different types of patents, suggesting the need for improvements to enhance model robustness and adaptability to diverse content types.
D.4 Evaluation on Alternative Prompts
To investigate the influence of different prompts, we have applied an alternative prompt to this task while the other settings remain unchanged: You are a patent expert. Given the following patent description, generate patent claims. Ensure the claims include all essential features, the language is unambiguous, the terminologies are used consistently, and features are interconnected and related accurately. Table 11 shows that all models perform better or maintain similar scores when using this revised prompt. Interestingly, the alternative prompt only brings marginal improvements to SaulLM-7B, Llama-3-8B, Llama-3-70B, and Mixtral-8$\times$7B, whereas Mistral-7B and GPT-4 are more sensitive to different prompts.
Appendix E Human Rating Strategies